Transform the general equation of circle : x2 + y2 + 4x - 2y -11 =0 to its standard form or center-radius form. After transforming , identify the center and the radius then sketch the graph in the Cartesian coordinate plane .

1
Expert's answer
2022-02-20T17:19:23-0500

Solution

For given equation

x2 + y2 + 4x - 2y - 11 = 0

x2 + 4x + y2 - 2y - 11 = 0

x2 + 4x + 4 + y2 - 2y + 1 - 11 – 5 = 0

(x + 2)2 + (y - 1)2 – 16 = 0

(x + 2)2 + (y - 1)2 = 42  

It is equation of the circle with center (xc,yc) = (-2, 1) and radius r = 4.